OpenStudy (anonymous):

sin 2x = _____ what is the best answer a. 2sinx + 2cos x b. 1/2 (cos(a-b) - cos(a+b)) c 2sinx cosx d 1-2sin^2 x

Parth (parthkohli):

Hint:\[\rm \sin(\alpha + \beta) = \sin \alpha \cos \beta + \sin \beta \cos \alpha\]You have to find sin(x + x).

OpenStudy (anonymous):

um ..what?

Parth (parthkohli):

Do you know what sin(x + y) is?

OpenStudy (anonymous):

Simplify for the answer.